OpenStudy (anonymous):

ok i dont just want an answer i want to know how, √(x)+2=x-4 haha i know the answer is 9 i dont understand how you get it. i understand squaring everything to get rid of the square root but i dont understand how u get 9 or what you do first! bahh!!!

OpenStudy (cathyangs):

ok...yes you square everything, but you first simplify the equation. srpt(x)=x-6 (you could figure it out from this if you know your squares) squaring it: x=x^2 -12x +36 then 0= x^2 -11x+36. You can solve this with the quadratic formula.